This the answer a got for part A of the problem I submitted which is 1B
1.Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: -2, where C is a constant and r is the distance in miles that the object is from the center of the earth.
1A) Solve the equation w = Cr-2 for r. I do have the answer, so could please answer the question 1B,

This is my question
1B. Suppose that an object is 150 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the earth

Answer by stanbon(75887)

1.Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: -2, where C is a constant and r is the distance in miles that the object is from the center of the earth.
1A) Solve the equation w = C*r^(-2) for r. I do have the answer, so could please answer the question 1B,
This is my question
1B. Suppose that an object is 150 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the earth
-----------
w = C*r^(-2)
150 = C*3963^-2
----
150*3963^2 = C
---
C = 2,355,805,350
================
Cheers,
stan H.